OpenStudy (anonymous):

verify that the given point is on the curve nd find the lines that are a) tangent and normal to the curve at the given point. x^2 -squareroot3xy + 2y^2 = 5 (sqrt3,2)

OpenStudy (anonymous):

is it \[x ^{2}-\sqrt{3x}y + 2y ^{2}=5 \] or \[x ^{2}-\sqrt{3xy} + 2y ^{2}=5 \] very important